OpenStudy (anonymous):

How would i take the derivative of this ? y=(sqrtx)e^(x^2)(x^2+8)^10

OpenStudy (anonymous):

\[(fgh)'=f'gh+fg'h+fgh'\]

OpenStudy (anonymous):

3(x-1)^2-4(x-1)+4use \[f(x)=\sqrt{x},f'(x)=\frac{1}{2\sqrt{x}}\]\[ g(x)=e^{x^2}, g'(x)=2xe^{x^2}\]\[h(x)=(x^2+8)^{10}, h'(x)=20x(x^2+8)^9\]

OpenStudy (anonymous):

okay thanks !